OpenStudy (anonymous):

a minor league baseball team plays 122 games a season. If the team won 14 more than three times as many games as they lost how many wins and losses did the team have?

ganeshie8 (ganeshie8):

say number of wins = W number of losses = L

ganeshie8 (ganeshie8):

W+L = 122

ganeshie8 (ganeshie8):

W = 3L + 14

ganeshie8 (ganeshie8):

two equations, two unknowns u can solve them ? :)

OpenStudy (anonymous):

wait huh?

ganeshie8 (ganeshie8):

W+L = 122 W = 3L + 14

ganeshie8 (ganeshie8):

these are two equations with two unknowns, u knw how to solve them ?

OpenStudy (anonymous):

nope not these ones

ganeshie8 (ganeshie8):

okay its easy simply put W value from second equation, in first equation : 3L+14+L = 122 4L + 14 = 122 subtract 14 both sides

ganeshie8 (ganeshie8):

4L = 108 divide by 4 L = 27 so, they have Lost 28 games ok

OpenStudy (anonymous):

it says its wrong I do my hw for college online n I types in 28 it says its wrong

ganeshie8 (ganeshie8):

there are two questions here : how many wins ? how many losses ?

ganeshie8 (ganeshie8):

27 is the answer for second question (losses)

OpenStudy (anonymous):

ok I got it thanks lol

ganeshie8 (ganeshie8):

im glad u got it :)